OpenStudy (anonymous):

PLS HELP?? During a science experiment, Faye recorded the temperature of a solution as it cooled in a controlled environment for 6 minutes. She graphed the change in temperature as shown below. If the solution continues to cool at the same rate, how many minutes will it take for the temperature to cool to 0°C? A. 24 B. 19 C. 14 D. 9

hartnn (hartnn):

can u find the equation of that line? by selecting any two points on that line.....

OpenStudy (anonymous):

Uhhh...ok so it would be the points (0, 40) and (6, 30) and m = y2- y1/x2 - x1, so m = 40 - 30 / 0 - 6 ?

OpenStudy (anonymous):

am i correct?

hartnn (hartnn):

yup,correct :) go ahead...

hartnn (hartnn):

yup,correct :) go ahead...

OpenStudy (anonymous):

ok so m = - 10/6 y - 30 = - 10/6 (x - 40) y - 30 = - 10/6 x - 40 + 40 + 40 so would the answer be y = - 10/6 x + 80 ?

OpenStudy (anonymous):

wait omg im so confused im sorry im dumb

hartnn (hartnn):

y - 40 = - 10/6 (x - 0) as the point (x1,y1) is (0,40) got this?

OpenStudy (anonymous):

yes ok then y - 40 = - 10/6 (x - 0) y - 40 = - 10/6 x - 0 + 0 + 0 y = - 10/6 x + 0 ????

hartnn (hartnn):

y - 40 = - 10/6 x - 0 adding 40 on both sides will give u y=-10/6x+40 ok?

OpenStudy (anonymous):

ohhh ok i was supposed to add 40 lol ok that makes alot more since

hartnn (hartnn):

or u can keep that as y - 40 = - 10/6 x now to get temperature 0, u put y=0 in that,to solve for x..... can u ?

OpenStudy (anonymous):

ahh im really confused on how to do this...wouldnt it just be the same thing if 0 - 40 = - 10/6 x 40 = - 10/6 x ??

hartnn (hartnn):

when u multiply by -1 -40*-1 = 40,thats correct but, - 10/6 x *(-1)=10/6 x so u should have 40=10/6 x now can u find x?

OpenStudy (anonymous):

sorry im actually really sorry lol omg im still really confused on how to find x ..

hartnn (hartnn):

but u clear till 40=10/6x right? so dividing by 10 u get 4=(1/6)x and multiplying by 6 gives x=6*4= ??

OpenStudy (anonymous):

24 yeah oh oh my god ok thank u soooooooo much wow it makes soooo much more since now. thank u!!!!!!!!!!

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

24